OpenStudy (anonymous):

Which of the following are solutions to the equation sinx cosx =-1/4? Check all that apply. A.7pi/12+npi B.11pi/12+npi C.7pi/12+npi/2 D.7pi/6+npi

OpenStudy (arindameducationusc):

what is n?

OpenStudy (anonymous):

I'm not sure, just part of the problem

OpenStudy (arindameducationusc):

Ok got it... wait.

OpenStudy (freckles):

use sin(2x)=2sin(x)cos(x) this is the double angle identity divide 2 on both sides you have 1/2*sin(2x)=sin(x)cos(x) replace sin(x)cos(x) with 1/2*sin(2x)

OpenStudy (arindameducationusc):

I agree with @freckles You will get x=pi/12 Then you can solve for n. Did you get it? Or you need more help

OpenStudy (anonymous):

I got it. The answer was A and B

OpenStudy (arindameducationusc):

Congrats @melisa1m